find the surface area of a cylinder with height 3m and radius 8.4m
pi = 3.14

Answer by checkley77(12844)

area=2*base+(circumferance*height)
area=2*pi*r^2+(2pi*r*3)
area=2*3.14*8.4^2+(2*3.14*8.4*3)
area=6.28*70.56+158.256
area=443.1168+158.256
area=601.37 m^2 ans.